OpenStudy (anonymous):

Can anyone plase help me? The general form of the equation of a circle is Ax2 + By2 + Cx + Dy + E = 0, where A = B 0. If the circle has a radius of 3 units and the center at (0, 4), which set of values of A, B, C, D, and E correspond to the circle?

OpenStudy (anonymous):

A = 0, B = 0, C = 2, D = 2, and E = 3 A = 1, B = 1, C = 8, D = 0, and E = 9 A = 1, B = 1, C = 0, D = -8, and E = 7 A = 1, B = 1, C = -8, D = 0, and E = 0 A = 1, B = 1, C = 8, D = 8, and E = 3

OpenStudy (janu16):

The other form of circle equation is (x-a)2 + (y-b)2 = r2. a and b is the coordinate of the center and r is radius. So the equation is x2+(y-4)2=9. Change it to the general form is x2+y2-8y+7=0. So A=1, B=1, C=0, D=-8 and E=7.

OpenStudy (janu16):

@McFly_47

OpenStudy (anonymous):

@Janu16 Thank you so much! I really appreciate your help

OpenStudy (janu16):

Yw